How to answer using MATLAB codes?

7. f(x) = Vx(x2 + 1), find f(x) and f(x). Simplify your final answers. (10pts) 8. y = sec? x + 3y tan x, solve the differ
Answer #1

Hi, as you didn't specify which ques to solve I am solving first two questions as per time limit, if you need further help or didn't understand anything in solution do comment I'll be happy to help you.

If you like my solution do leave positive comment and a thumbs up.

I have tested all the code in online compiler octave

Code:

1)code for f'(x) or first order directive

syms x
f= sqrt(x)*(x^2 + 1)
diff(f)

Code for f'''(x) or third order derivative

syms x
f= sqrt(x)*(x^2 + 1)
diff(f,3)

2)

syms y(x)  
ode = diff(y,x) == sec(x)^2 + 3*y*tan(x) 
#ode = diff(y,t) == t*y
ySol(x) = dsolve(ode)